OpenStudy (anonymous):

Limit of (x^2 - 4x + 3) / cube root of (x) - 1 as X approaches 1

OpenStudy (shubhamsrg):

x^2 - 4x + 3 = (x-1)(x-3) now x-1 = cuberoot(x) ^3 - (1)^3 apply a^3 - b^3 formula . This should help you.